a. In the country of Brazistan, two types of cigarette are marketed: local brand cigarettes, which sell at the price before taxes of $0.50 per packet of 20, and luxury international brand cigarettes, which sell at $1.50 per packet of 20. Without taxes, it is estimated that annual sales would be 2,000 million packets of local cigarettes and 250 million packets of international brand cigarettes. Brazistan has a retail sales tax of 20%. If the production of cigarettes is characterized by constant costs, and if the own-price elasticity of demand for local brand cigarettes is –0.8 and for international brand cigarettes is –1.1, then estimate the annual quantity, tax revenue and excess burden of the tax on each brand of cigarette. Also express the excess burden as a share of the revenues.
Answer #1

Local brand cigrettes

Price Elasticity= % change in quantity/% change in price

% change in quantity= 20*-0.8= -16%

Quantity falls 16% from 2000.

New quantity= 1680

Tax= 0.20*$0.50= $ 0.10

Tax revenue=Tax* New quantity = $0.10*1680= $168

Excess burden of tax= Deadweight loss= 1/2*tax*change in quantity= 1/2*0.10*2000-1680= $16

Excess burden as a share of revenue= $16/168= 0.095 or 9.5%

International brand cigrettes

Price Elasticity= % change in quantity/% change in price

% change in quantity= 20*-1.1= -22%

Quantity falls 22% from 250

New quantity= 195

Tax= 0.20*$1.5= $0.3

Tax revenue=Tax* New quantity = $0.3*195= $58.5

Excess burden of tax= Deadweight loss= 1/2*tax*change in quantity= 1/2*0.3*250-195= $8.25

Excess burden as a share of revenue= $8.25/$58.5= 0.141 or 14.1%